OpenStudy (anonymous):

Find the length of the radius of the given circle. (x - 5)2 + (y + 3)2 = 25 5 12.5 25

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(x-a)^2+(y-b)^2=r^2 where (a, b) is the center r radius

OpenStudy (alexandervonhumboldt2):

sqrt(r^2)

OpenStudy (alexandervonhumboldt2):

r^2=25

OpenStudy (alexandervonhumboldt2):

sqrt25=5

OpenStudy (anonymous):

thnx